OpenStudy (anonymous):

Select the coordinates of two points that are on the line x = 5.

OpenStudy (anonymous):

Do you have answer options?

OpenStudy (anonymous):

(5, 0) and (0, 5) (5, 0) and (5, -2) (5, 0) and (0, 2) (0, 5) and (0, 2)

OpenStudy (anonymous):

B. (5,0) and (5,-2)

OpenStudy (anonymous):

can i ask more questions

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4). (2 points) y = x + 8 y = x + 3 y = - x + 8 y = - x + 3

OpenStudy (anonymous):

To make it a little easier think of this a coordinate has location (x, y), you are looking for two points in which x = 5, which is a vertical line... so just find the two points that has a 5 in the first spot.

OpenStudy (anonymous):

c right

OpenStudy (anonymous):

??

OpenStudy (anonymous):

It's parallel, so it goes in the same direction, meaning it has the same slope. Are there fractions in front of the x's that didn't get copied?